OpenStudy (anonymous):

Oil is delivered to a refinery between the hours of midnight (t = 0) and 5 a.m. (t = 5). Selected values of the rate of delivery are shown in the table below, with t measured in hours past midnight, and R(t) measured in gallons per hour. Use 4 trapezoids to estimate the total amount of oil to the nearest gallon that is delivered between midnight and 5 a.m.

OpenStudy (anonymous):

@terenzreignz @jim_thompson5910 @zepdrix This is fairly easy just want to confirm my answer. I got 1075

jimthompson5910 (jim_thompson5910):

you forgot to post the table

OpenStudy (anonymous):

https://gyazo.com/9252eae33a7f37da2771d882787b4a8c

OpenStudy (anonymous):

My apologies

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

I was reading another problem, my bad

jimthompson5910 (jim_thompson5910):

I'm hitting a stumbling block. It seems like the "4 trapezoids" should be "5 trapezoids" 4 doesn't work because you have the right edges that don't land on given table values

jimthompson5910 (jim_thompson5910):

I guess we could leave off the last column of the table since the 5 am cut off happens there

OpenStudy (anonymous):

What? I was able to do it with 4 no problem unless you see something I don't

jimthompson5910 (jim_thompson5910):

oh nvm, I mistakenly put 3 in there

jimthompson5910 (jim_thompson5910):

I'm getting something larger than 1075

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

I'll write out what U did.

OpenStudy (anonymous):

I*

jimthompson5910 (jim_thompson5910):

here is what the trapezoids look like

OpenStudy (anonymous):

\[(\frac{ 100+250 }{ 2 })+(\frac{ 250+200 }{ 2 })+( 200+350)+(\frac{ 350+100 }{ 2 } )\]

jimthompson5910 (jim_thompson5910):

yep and that leads to the answer I'm getting, which is 1175

OpenStudy (anonymous):

Probably miscalculated. Let me give it another go

OpenStudy (anonymous):

You are right.